OpenStudy (anonymous):

Find the function y=y(x) (for x>0) which satisfies the separable differential equation dy/dx=5+12x/xy^2; x>0 with the initial condition y(1)=6 solve for y.

myininaya (myininaya):

\[\frac{dy}{dx}=5+\frac{12x}{xy^2}\] I don't think it can be done with separation of variables

OpenStudy (anonymous):

how do we solve then?

myininaya (myininaya):

well we can cancel an x in the second term so we have \[y'=5+\frac{12}{y^2}\] thinking...

myininaya (myininaya):

\[y^2y'=5y^2+12 => y^2y'-5y^2=12\] hmm i can't think of how to do this one

OpenStudy (anonymous):

im not sure how to deal with it without the seperation of variables

myininaya (myininaya):

did you write the problem down correctly because you can't use separation on variables on this one you could do it for this one: \[\frac{dy}{dx}=\frac{5+12x}{xy^2}\] but you wrote dy/dx=5+12x/xy^2 instead of dy/dx=(5+12x)/xy^2

OpenStudy (anonymous):

thats my mistake you are right

myininaya (myininaya):

ok goodie we can do this one then

myininaya (myininaya):

i will start by multiplying both sides by y^2 \[y^2 \frac{dy}{dx}=\frac{5+12x}{x}\] now i will multiply dx on both sides \[y^2 dy=\frac{5+12x}{x} dx\] now integrate both sides \[\frac{y^{2+1}}{2+1}+C_1=5\ln|x|+12x+C_2\]

myininaya (myininaya):

\[\frac{y^3}{3}=5\ln|x|+12x+C\]

myininaya (myininaya):

since \[C_2-C_1=constant \]

OpenStudy (anonymous):

solve for c?

myininaya (myininaya):

\[y^3=15\ln|x|+36x+C\]

myininaya (myininaya):

yes apply initial condition to find C

OpenStudy (anonymous):

C=180

OpenStudy (anonymous):

so if c=180 then \[y=\sqrt[3]{15\ln(x)+36x+180}\]

myininaya (myininaya):

6^3=36+c 36(6)=36+C 216=36+C 180=C yes looks awesome :)

OpenStudy (anonymous):

thank you